how do i write and equation of the line through points (2,4) and (3,-3)?
Answer by nerdybill(7384)

how do i write and equation of the line through points (2,4) and (3,-3)?
first, find slope between the two points by applying:
m = (y2-y1)/(x2-x1)
m = (-3-4)/(3-2)
m = -7/1
m = -7
.
Using the slope (m) and either of the two points (2,4) plug it into the "point-slope" form:
y - y1 = m(x - x1)
y - 4 = -7(x - 2)
y - 4 = -7x + 14
y = -7x + 18 (this is the answer in "slope-intercept" form)